OpenStudy (anonymous):

Using the equation below as a model, fill in numbers in the place of a and b to create a rational equation that has an extraneous solution.

OpenStudy (anonymous):

\[\frac{ x+a }{ ax}=\frac{ b }{ x }\]

OpenStudy (anonymous):

Part 1. Show all work to solve for x in the equation and check the solution. Part 2. Explain how to identify the extraneous solution and what it means.

OpenStudy (anonymous):

i got this is it right? (x+a)/ax=b/x x/ax+a/ax=b/x 1/a+1/x=b/x x+1=b x=b-1

OpenStudy (anonymous):

@ninjasandtigers

OpenStudy (anonymous):

yeah that's right

OpenStudy (anonymous):

and how would i check if that is an extraneous solution can you show me the steps?

OpenStudy (anonymous):

@satellite73